Organic Oranges Bag Company has the following figures for August 2020: Bags sold: 40,000 Price per bag: $10 Variable cost per bag: $4 Fixed cost: $120,000 (1) Calculate the breakeven number in units for August. (2) How many bags does the firm need to sell in September to make a target profit of $150,000? Please show all work for the problem.
Answer #1

Contribution margin=Sales-Variable cost  

=(10-4)=$6 per bag

a.Breakeven=Fixed expenses/Contribution margin

=120,000/6

=20,000 bags

b.Target contribution margin=Fixed cost+Target profits

=(120,000+150,000)=$270,000

Hence bags to be sold=Target contribution margin/contribution margin per unit

=270,000/6

=45,000 bags